OpenStudy (anonymous):

Write an exponential form with prime base: (8^3*4^5)^2/ 32^5

OpenStudy (welshfella):

Convert to powers of 2 8 = 2^3 , 4 = 2^2 , 32 = 2^5

OpenStudy (welshfella):

so for example 8^3 = (2^3)^3 = 2^9

OpenStudy (welshfella):

finally simplify